OpenStudy (anonymous):

HELP!!! Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). y = one thirdx − 3 y = one thirdx + 17 y = −3x − 3 y = −3x + 17

HanAkoSolo (jamierox4ev3r):

3x+y=7 -3x -3x y=-3x+7

OpenStudy (anonymous):

thx

HanAkoSolo (jamierox4ev3r):

wait but if it is perpendicular you must remember to use the inverse, which would have a slope of 1/3x okay?

OpenStudy (anonymous):

k